OpenStudy (anonymous):

write an equation for a line in slope-intercept form passing through (3,11) and parallel to y = 7. A. y = 11 B. x = 11 C. x = 3 D. y = 3

sam (.sam.):

\[\huge y-y _{1}=m(x-x _{1})\]

OpenStudy (hoblos):

parallel to y = 7 then it is horizontal line of slope 0 y=11

Directrix (directrix):

A. y = 11

Directrix (directrix):

y - 11 = 0(x-3) y = 11